A certain airplane has two independent alternators to provide electrical power. The probability that a given alternator will fail on a 1-hour flight is .02. What is the probability that (a) both will fail? (b)Neither will fail? (c) One or the other will fail? Show all steps.
Answer by sudhanshu_kmr(1152)


probability of fail = 0.02 = 2/100
probability of not fail = 1 / 2/100 = 98/100

(A) both will fail = 2/100 * 2/100 = 4/10000


(b) neither will fail = 98/100 * 98/100


(c) one or other will fail = 2/100 * 98/100 + 98/100 * 2/100
= 392/10000